# WIMP/FIMP dark matter and primordial black holes with memory burden effect

### v1

@@ Abstract
The lifetime of primordial black holes (PBHs), which were formed in the early universe, can be extended by the memory burden effect. Light PBHs may be in existence today and become a candidate for dark matter (DM). We assume that DM is made of thermally produced weakly interacting massive particles (WIMPs), WIMPs produced via Hawking radiation of PBHs, and survived PBHs via the memory burden effect. In addition, feebly interacting massive particles (FIMPs) are considered as an alternative to WIMPs. Various similar studies have been conducted either with or without considering the memory burden effect. In this study, we simultaneously account for thermally produced WIMPs or FIMPs and the memory burden effect of PBHs to explain the relic abundance of DM. We show that this abundance is highly sensitive to the memory burden effect in PBHs.
@@ 1 Introduction
The existence and production mechanism of dark matter (DM) have been a longstanding mystery in cosmology and particle physics [1]. The traditional and most studied thermal DM production mechanism is freeze-out (FO) [2]. In this mechanism, DM particles $\chi$ reach thermal equilibrium with thermal bath particles $f$ in the early universe through annihilation and pair production $\chi\chi\leftrightarrow\bar{f}f$. Subsequently, the temperature of bath particles drops below the mass of DM particles and $\Gamma\lesssim H$ is satisfied, where $\Gamma$ is the reaction rate of annihilation $\chi\chi\rightarrow\bar{f}f$ and $H$ denotes the Hubble parameter. Then, the comoving number density of DM particles remains constant. As a typical weak-scale cross-section leads to the observed relic abundance of DM, the FO scenario is usually discussed in the context of weakly interacting massive particles (WIMPs). Although WIMPs remain a leading candidate explanation for DM, they elude detection in numerous experiments. Thus, other candidates have been proposed, such as feebly interacting massive particles (FIMPs) [3, 4] based on the opposite idea of WIMPs. If the coupling of DM particles $\chi$ to bath particles is very weak, DM particles never reach thermal equilibrium with the thermal bath. Such a DM particle is called a FIMP. Assuming that the initial density of FIMPs is negligible, the feeble interaction with the thermal bath leads to FIMP production during the expansion of the universe. This DM production scheme is called the freeze-in (FI) mechanism and is applicable to different scenarios such as decay $f_{1}\rightarrow f_{2}\chi$, scattering $f_{1}f_{2}\rightarrow f_{3}\chi$, and pair production $f_{1}f_{2}\rightarrow\chi\chi$.
@@
Primordial black holes (PBHs), which were formed in the early universe, are also feasible candidates for explaining DM [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. PBHs emit particles through Hawking radiation [17], which reduces their mass. Therefore, PBHs have a finite lifetime. However, if this lifetime is longer than the age of the universe, PBHs may exist until the present day without being completely evaporated.
@@
PBHs that formed with a mass of less than approximately $10^{15}$ g were believed not to survive to the present day. Nevertheless, it has recently been discovered that the lifetime of PBHs can be extended by the memory burden effect [18, 19, 20, 21, 22]. The quantum corrections for PBHs due to the memory burden effect are being extensively studied to explain the relic abundance of PBH [23, 26, 24, 25], axion DM [27], gravitational waves [28, 29, 26], baryon asymmetry [32, 30, 31], cosmic neutrinos [33, 34, 35], compact stars [36], and phenomena beyond the Standard Model of particle physics [37].
@@
We assume that DM is made of both WIMPs (or FIMPs) and survived PBHs. As Hawking radiation is induced by gravity, PBHs evaporate into all particle species. WIMPs (FIMPs) are also produced by Hawking radiation. Thus, we should treat the following three DM components:
  • Thermally produced WIMPs (or FIMPs),
  • WIMPs (or FIMPs) produced from Hawking radiation of PBHs, and
  • Survived PBHs.
@@
Similar studies have neglected the memory burden effect [39, 38, 40, 41, 42, 43, 44, 45, 46, 47]. Considering the memory burden effect, a study attempted to explain the relic abundance of DM via Hawking radiation of PBHs and survived PBHs [23]. However, this study did not account for DM related to thermally produced particles.
@@
We simultaneously consider thermally produced WIMPs/FIMPs and the memory burden effect of PBHs to explain the relic abundance of DM. If the effect of thermal production of WIMPs or FIMPs is much smaller than that of the particle production by PBHs via Hawking radiation, we can omit the contribution of the thermally produced WIMPs or FIMPs. In this case, the results are the same as those reported in [23]. Therefore, we focus on the case that the effect of thermal production of WIMPs or FIMPs is much larger than the effect of the particle production by PBHs via Hawking radiation.
@@
The remainder of this paper is organized as follows. Section 2 provides a brief review of PBHs with the memory burden effect. In Sec. 3, we assess the combined effects of PBHs and thermally produced WIMPs or FIMPs on the relic abundance of DM. Finally, Sec. 4 provides a summary of the study and findings.
@@ 2.1 Time evolution
$\Gamma_{\rm BK}^{0}=\frac{3\epsilon M_{\rm P}^{4}}{M_{\rm in}^{3}}.$ (5)
@@
$\tau^{0}=t_{\rm ev}^{0}-t_{\rm in}=\frac{1}{\Gamma_{\rm BK}^{0}}\simeq 2.4\times 10^{-28}\left(\frac{M_{\rm in}}{1{\rm g}}\right)^{3}\ {\rm s},$ (6)
@@
is the black hole entropy. Parameter $k$ characterizes the efficiency of the back reaction effect, with $k=0$ representing the case of Hawking radiation without the memory burden effect.
We assume that the semiclassical era is valid until the PBH mass reaches
@@
with $10<q<$ 1. From Eq. (4), we define the final time for the semiclassical era as
@@
for $k\neq 0$ and $10<q<$ 1. We estimate the evaporation time by solving $M_{\rm BH}(t_{\rm ev}^{k})=0$ as follows [20, 23, 32]:
$t_{\rm ev}^{k}=t_{q}+\frac{1}{\Gamma_{\rm BK}^{k}}\simeq(1-q^{3})\frac{M_{\rm in}^{3}}{3\epsilon M_{\rm P}^{4}}+\frac{1}{2^{k}(3+2k)\epsilon M_{\rm P}}\left(\frac{qM_{\rm in}}{M_{P}}\right)^{3+2k}.$ (13)
Approximation $t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BH}^{k}$ is adequate in some cases.
@@ 2.2 Initial density
The initial density of a PBH can be described by the following ratio:
@@
where $\rho_{\rm BH}^{\rm in}$ and $\rho_{\rm r}^{\rm in}=\pi^{2}g_{\ast}(T_{\rm in})T_{\rm in}^{4}/30$ are the initial energy densities of the PBH and radiation, respectively. We assume that the PBH formed during a radiation-dominated era in the universe. Hence, the radiation energy density varies as follows:
@@
As a PBH behaves as matter, its energy density evolves as follows:
@@
where $a$ is a scale factor, and instant mass reduction is assumed as $M_{\rm in}\rightarrow qM_{\rm in}$ at $a=a_{\rm q}$. From $\rho_{\rm r}=\rho_{\rm BH}$, the radiation–PBH equivalent may occur at $a_{\rm BH}=a_{\rm in}/(q\beta)$. A PBH-dominant era is possible if $a_{\rm BH}<a_{\rm ev}$, which corresponds to
@@
with $\gamma\simeq 0.2$. We assume that the initial PBH mass is related to the horizon size at PBH formation as follows:
@@
In addition, we use relations $H=1/(2t)$, $a\propto t^{1/2}$, and $H_{\rm in}^{2}=\rho_{\rm r}^{\rm in}/(3M_{\rm P}^{2})$ for a radiation-dominated universe, and $t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BK}^{k}$ using Eq. (12).
As addressed in [23], $\beta<\beta_{c}$ establishes a case of interest because we assume that PBHs evaporate completely after the present time due to the memory burden effect and survived PBHs until the present day are candidates of DM (no early matter-dominated era is considered). Hereafter, we require relation $\beta<\beta_{c}$ to hold.
@@ 2.3 Survived PBH density
We assume that PBHs evaporate completely after the present time due to the memory burden effect. Thus, the density of survived PBHs is obtained as follows [48]:
@@ 2.4 Particle production
The number of particle species, $i$, from PBH evaporation is controlled by the relation between particle mass $m_{i}$ and the PBH formation temperature given by
@@
If $m_{i}<T_{\rm in}$, particle emission from a PBH occurs from $t_{\rm in}$ to $t_{\rm ev}$. If $m_{i}>T_{\rm in}$, particle emission starts from time $t_{i}$, where $m_{i}=T_{\rm BH}(t_{i})$. The relic abundance of emitted particle DM is given by Eq. (3.33) in [23]:

### v3

@@ Abstract
The lifetime of primordial black holes (PBHs), which formed in the early universe, can be extended by the memory burden effect. Light PBHs may exist today and be candidates for dark matter (DM). We assume that DM is made of thermally produced weakly interacting massive particles (WIMPs), WIMPs produced via the Hawking radiation of PBHs, and PBHs that survived Hawking evaporation via the memory burden effect. Feebly interacting massive particles (FIMPs) are alternatives to WIMPs. This paper shows that a small memory burden is preferable if the thermal production of WIMPs or FIMPs is much larger than the effect of PBH particle production via Hawking radiation. In addition, we show that the lower limit of the DM mass, called the warm DM (WDM) constraint, decreases with the memory burden effect of PBHs. Results suggest that the WDM constraint is more effective for FIMPs than for WIMPs.
{history}
@@ 1 Introduction
The existence and production mechanism of dark matter (DM) are a long-standing mystery in cosmology and particle physics [1]. The traditional and most studied thermal DM production mechanism is freeze out (FO) [2]. In this mechanism, DM particles, $\chi$, reached thermal equilibrium with thermal bath particles, $f$, in the early universe through annihilation and pair production, $\chi\chi\leftrightarrow\bar{f}f$. Subsequently, the temperature of the bath particles drops below the mass of the DM particles and $\Gamma\lesssim H$ holds, where $\Gamma$ is the reaction rate of annihilation, $\chi\chi\rightarrow\bar{f}f$, and $H$ is the Hubble parameter. Then, the comoving number density of the DM particles remains constant. A typical weak-scale cross section leads to an observed relic abundance of DM, so the FO scenario is usually discussed in the context of weakly interacting massive particles (WIMPs). Although WIMPs are a leading candidate explanation for DM, they cannot be detected in numerous experiments. Thus, feebly interacting massive particles (FIMPs) [3, 4] have been proposed based on an idea opposite that of WIMPs. FIMPs are DM particles that never reached thermal equilibrium with thermal bath particles because of very weak coupling between them. Assuming a negligible initial density of FIMPs, the feeble interaction of DM particles with thermal bath particles leads to FIMP production during the expansion of the universe. This DM production scheme, called the freeze-in (FI) mechanism, is applicable to scenarios such as decay ($f_{1}\rightarrow f_{2}\chi$), scattering ($f_{1}f_{2}\rightarrow f_{3}\chi$), and pair production ($f_{1}f_{2}\rightarrow\chi\chi$).
@@
Primordial black holes (PBHs), which formed in the early universe, are also feasible candidates for explaining DM [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. PBHs emit particles through Hawking radiation [18], which reduces their mass. Therefore, PBHs have a finite lifetime. However, if this lifetime is longer than the age of the universe, then some PBHs may not completely evaporate and may exist today.
@@
PBHs that formed with a mass of less than approximately $10^{15}$ g were believed not to have survived to the present day. However, recent findings suggest that the lifetime of PBHs can be extended by the memory burden effect [19, 20, 21, 22, 23, 24]. Quantum corrections for PBHs due to the memory burden effect are being studied extensively to explain the relic abundance of PBHs [25, 28, 26, 27], gravitational waves (GWs) [29, 30, 31, 32, 33, 34, 28, 35, 36, 37], axion DM [38], baryon asymmetry [41, 39, 40], cosmic neutrinos [42, 43, 44, 45], compact stars [46], and phenomena beyond those in the standard model (SM) of particle physics [47].
@@
We assume that DM consists of WIMPs/FIMPs and survived PBHs. As Hawking radiation is induced by gravity, PBHs evaporate into all particle species. WIMPs/FIMPs are also produced by Hawking radiation. Thus, we consider the following DM components:
  • thermally produced WIMPs/FIMPs
  • WIMPs/FIMPs produced by the Hawking radiation of PBHs
  • survived PBHs
@@
Similar studies neglect the memory burden effect [49, 48, 50, 51, 52, 53, 54, 55, 56, 57]. Considering the memory burden effect, researchers attempted to explain the relic abundance of DM via the Hawking radiation of PBHs and survived PBHs [25]. However, they did not account for DM related to thermally produced particles.
@@
We simultaneously consider thermally produced WIMPs/FIMPs and the memory burden effect of PBHs to explain the relic abundance of DM. We can omit the contribution of thermally produced WIMPs/FIMPs to the relic abundance of DM if the effect of their thermal production is much smaller than that of particle production by PBHs via Hawking radiation. In this case, the results are the same as those previously reported [25].
According to Ref. [25], we discuss a case where the semiclassical phase of Hawking evaporation ends before big bang nucleosynthesis (BBN) and the PBHs evaporate completely after the present time due to the memory burden effect (the so-called early-matter-dominated era does not exist). To examine a setup different from those in previous studies, we focus on a case where the effect of the thermal production of WIMPs/FIMPs considerably exceeds that of PBH particle production via Hawking radiation.
@@
The remainder of this paper is organized as follows. Section 2 provides a brief review of PBHs with the memory burden effect. In Section 3, we discuss the constraints on PBHs from BBN, cosmic microwave background radiation (CMB), GWs, and warm DM (WDM).We show that the lower limit of the DM mass (the WDM constraint, which is related to structure formation in the universe) decreases with the memory burden effect of PBHs. In Section 4, we assess the combined effects of PBHs and thermally produced WIMPs/FIMPs on the relic abundance of DM. Finally, Section 5 provides a summary of this study and its findings.
@@ 2.1 Time evolution
$\Gamma_{\rm BH}^{0}=\frac{3\epsilon M_{\rm P}^{4}}{M_{\rm in}^{3}}.$ (5)
@@
$\tau^{0}=t_{\rm ev}^{0}-t_{\rm in}=\frac{1}{\Gamma_{\rm BH}^{0}}\simeq 2.4\times 10^{-28}\left(\frac{M_{\rm in}}{1{\rm g}}\right)^{3}\ {\rm s},$ (6)
@@
is the black hole entropy. The parameter $k$ characterizes the efficiency of the back reaction effect, with $k=0$ indicating Hawking radiation without the memory burden effect.
The semiclassical era is assumed valid until the PBH mass reaches
@@
where $0<q<1$. From Eq. (4), we define the final time for the semiclassical era as
@@
at $k\neq 0$ and $0<q<1$. We estimate the evaporation time by solving $M_{\rm BH}(t_{\rm ev}^{k})=0$ as [21, 25, 41]
$t_{\rm ev}^{k}$ $=$ $t_{q}+\frac{1}{\Gamma_{\rm BH}^{k}}\simeq(1-q^{3})\frac{M_{\rm in}^{3}}{3\epsilon M_{\rm P}^{4}}$ (13)
$+\frac{1}{2^{k}(3+2k)\epsilon M_{\rm P}}\left(\frac{qM_{\rm in}}{M_{P}}\right)^{3+2k}.$
Approximation ($t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BH}^{k}$) is adequate in some cases.
@@ 2.2 Initial density
The initial density of a PBH is
@@
where $\rho_{\rm BH}^{\rm in}$ and $\rho_{\rm r}^{\rm in}=\pi^{2}g_{\ast}(T_{\rm in})T_{\rm in}^{4}/30$ are the initial energy densities of the PBH and radiation, respectively. We assume that the PBH formed during a radiation-dominated era in the universe. Hence, the radiation energy density varies as
@@
As a PBH behaves as matter, its energy density evolves as
@@
where $a$ is a scale factor and instant mass reduction is assumed as $M_{\rm in}\rightarrow qM_{\rm in}$ at $a=a_{\rm q}$. According to $\rho_{\rm r}=\rho_{\rm BH}$, the radiation–PBH equivalent may occur at $a_{\rm BH}=a_{\rm in}/(q\beta)$. A PBH-dominant era is possible if $a_{\rm BH}<a_{\rm ev}$, which corresponds to
@@
where $\gamma\simeq 0.2$. We assume that the initial PBH mass is related to the horizon size during PBH formation as
@@
In addition, we use relations $H=1/(2t)$, $a\propto t^{1/2}$, and $H_{\rm in}^{2}=\rho_{\rm r}^{\rm in}/(3M_{\rm P}^{2})$ for a radiation-dominated universe and set $t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BK}^{k}$ using Eq. (12).
As addressed in Ref. [25], $\beta<\beta_{\rm c}$ is a case of interest because we assume that PBHs will evaporate completely after the present time due to the memory burden effect and PBHs that have survived until the present day are DM candidates (no early-matter-dominated era is considered). Hereinafter, $\beta<\beta_{\rm c}$ should hold.
@@ 2.3 Survived PBH density
We assume that PBHs will evaporate completely after the present time due to the memory burden effect. Thus, the density of survived PBHs is [58]
@@ 2.4 Particle production
The number of particle species, $i$, from PBH evaporation is controlled by the relation between the particle mass, $m_{i}$, and the PBH formation temperature, which is
@@
If $m_{i}<T_{\rm in}$, then particle emission from a PBH occurs from $t_{\rm in}$ to $t_{\rm ev}$. If $m_{i}>T_{\rm in}$, then particle emission starts from time $t_{i}$, where $m_{i}=T_{\rm BH}(t_{i})$. The relic abundance of emitted particle DM is given by Eq. (3.33) in [25].
@@
3 Constraints
3.1 BBN
We focus on a case where the semiclassical phase of PBH evaporation ends before BBN. The initial mass of the PBH, which is partially evaporated until the final time for the semiclassical era, $t_{q}$, can be obtained from Eqs. (5), (6), and (10) as
$M_{\rm in,t_{q}}=1.61\times 10^{9}\left(\frac{1}{1-q^{3}}\right)^{1/3}\left(\frac{t_{q}}{1{\rm s}}\right)^{1/3}~{\rm g},$ (26)
where we assume $t_{\rm in}\ll t_{\rm ev}^{0}$.
In this study, we set the time of BBN as $t_{\rm BBN}=1$ s for simplicity. We obtain
$M_{\rm in,t_{\rm BBN}}=1.61\times 10^{9}\left(\frac{1}{1-q^{3}}\right)^{1/3}~{\rm g}$ (27)
at $t_{q}=t_{\rm BBN}$. If $M_{\rm in}>M_{\rm in,t_{\rm BBN}}$, then the semiclassical evaporation of PBHs happens after BBN. Because $1\leq(1-q^{3})^{-1/3}\leq 1.055$ at $0<q\leq 0.9$, $M_{\rm in,t_{\rm BBN}}$ is almost independent of $q$ in this study. We use $M_{\rm in}\leq M_{\rm in,t_{\rm BBN}}=1\times 10^{9}$ g as the BBN constraint, which is the same as that without the memory burden effect [7, 59, 60, 61, 49].
3.2 CMB
Constraints on inflation impose a lower limit on the PBH mass. Given that the Hubble parameter at a time is less than or equal to the Hubble parameter during inflation, we obtain the lower limit of the initial mass of the PBH as $M_{\rm in}\gtrsim 0.1$ g [49, 7]. Because this constraint comes from PBH generation, it is independent of the memory burden effect. Although this constraint depends on the details of gravitational collapse and the inflation scenario, we regard $M_{\rm in}\gtrsim 0.1$ g as the CMB constraint by assuming a standard slow-roll scenario for simplicity [7].
3.3 GWs
PBHs are potential sources of GWs [62, 63, 64, 65, 66, 67, 68, 69, 70, 72, 71, 73, 74, 75, 76, 77, 78, 79, 80]. PBH formation is accompanied by GWs, and primordial curvature perturbations can induce GWs in the final stage of inflation, immediately after horizon re-entry.
GWs can be induced by sources other than primordial fluctuations [81, 82, 83]. PBHs are randomly distributed in space according to approximate Poisson statistics immediately after formation. Although PBH gas behaves similarly to pressureless dust, its spatially inhomogeneous distribution causes isocurvature density fluctuations. At a sufficiently large initial fraction of PBHs in the early-matter-dominated (PBH-dominated) era ($\beta>\beta_{\rm c}$ case), the initial isocurvature fluctuations become curvature perturbations. These curvature perturbations then act as sources of secondary GWs [81, 82, 85, 83, 86, 87, 84, 88, 89]. These induced GWs can be detected via future GW observations, such as those obtained using Hanford–Livingstone–Virgo–KAGRA [90, 91, 92, 93], the Einstein Telescope [94, 95], Cosmic Explorer [96], Laser Interferometer Space Antenna [97], Deci-Hertz Interferometer Gravitational-Wave Observatory [98, 99], and Big Bang Observer [100, 101].
GWs induced by PBHs with the memory burden effect have been extensively studied recently [29, 30, 31, 32, 33, 34, 28, 35, 36, 37]. Findings suggest that the GW spectrum amplitude is enhanced by the longer early PBH-dominated phase caused by the memory burden effect, assuming the existence of the early-matter-dominated era ($\beta>\beta_{\rm c}$).
In this study, PBHs are assumed to evaporate completely after the present time due to the memory burden effect and PBHs that have survived until the present day are regarded as DM candidates ($\beta<\beta_{\rm c}$ should hold). We assume that the early-matter-dominated era did not exist and PBHs never dominated the energy density in the history of the universe. In addition, we focus on a case where the effect of the thermal production of WIMPs/FIMPs substantially exceeds that of particle production by PBHs. In this case, GWs from PBHs almost behave as if the memory burden effect is nonexistent, so a small GW amplitude is expected (see $q=0$ curve in Fig. 3 in Ref. [29] for example).
Figure 1: WDM constraint with memory burden effect of PBHs. Left panel: lower limit of WDM mass vs. initial mass of PBH. The left (right) gray region is excluded from the CMB (BBN) constraint. Right panel: $q$ dependence on lower limit of $m_{\chi}$.
3.4 WDM
DM should be cold enough to avoid erasing small-scale structures via free streaming. A constraint on WDM or noncold DM is obtained via the study of the Lyman-$\alpha$ forest [102, 103, 104, 105, 49, 106].
In cases with negligible DM interaction with any particles in the thermal bath, DM particles can only lose energy (or momentum) by redshifting. The present redshifted momentum of DM, $p(t_{0})$, is [48]
$p(t_{0})$ $=$ $\frac{a(t_{\rm ev})}{a(t_{0})}p(t_{\rm ev})$ (28)
$=$ $\frac{a(t_{\rm ev})}{a(t_{\rm eq})}\frac{a(t_{\rm eq})}{a(t_{0})}p(t_{\rm ev})$
$=$ $\frac{a(t_{\rm ev})}{a(t_{\rm eq})}\frac{\Omega_{\rm r}}{\Omega_{\rm m}}p(t_{\rm ev}),$
where $\Omega_{\rm r}$ and $\Omega_{\rm m}$ are the density parameters of radiation and matter, respectively. Using $\rho_{\rm r}\propto a^{-4}$, we obtain
$\frac{a(t_{\rm ev})}{a(t_{\rm eq})}=\left(\frac{\rho_{\rm r}(T_{\rm eq})}{\rho_{\rm r}(T_{\rm ev})}\right)^{1/4}.$ (29)
We can use the relations
$\rho_{\rm r}(T_{\rm eq})$ $=$ $\rho_{\rm c}\left(\frac{a(T_{0})}{a(T_{\rm eq})}\right)^{3}=\rho_{\rm c}a(T_{\rm eq})^{-3},$
$\rho_{\rm r}(T_{\rm ev})$ $=$ $3M_{\rm P}^{2}H^{2}=\frac{3M_{\rm P}^{2}}{(2t_{\rm ev})^{2}}\simeq\frac{3M_{\rm P}^{2}}{4\tau^{2}},$
$a(T_{\rm eq})$ $=$ $a(t_{\rm eq})=\frac{\Omega_{\rm r}}{\Omega_{\rm m}},$ (30)
and $p(t_{\rm ev})\simeq\langle E\rangle\simeq 6T_{\rm in}$ for light DM ($m_{\rm DM}\ll T_{\rm in}$), where $\rho_{\rm c}=3H_{0}^{2}/(8\pi G)=1.05h^{2}\times 10^{-5}\ {\rm GeV\ cm^{-3}}$ is the critical density of the universe and $\tau$ is the PBH lifetime.
Without the memory burden effect, we set $\tau=\tau^{0}$ under the assumption of $t_{\rm in}\ll t^{0}_{\rm ev}$. In this case, the momentum of a DM particle today is
$p(t_{0})=1.53\times 10^{-12}\left(\frac{M_{\rm in}}{M_{\rm P}}\right)^{1/2}~{\rm GeV}$ (31)
at $g_{*}(T_{\rm BH})=106.75$, $\Omega_{\rm r}=5.38\times 10^{-5}$, and $\Omega_{\rm m}=0.315$ [107]. Assuming the WDM is heavier than 3.5 keV [108], the upper limit of the DM velocity is approximately $v_{\rm DM}\lesssim 1.8\times 10^{-8}$ [109]. From $p(t_{0})=m_{\rm DM}v_{\rm DM}$, we obtain the following WDM constraint on the DM mass and the initial mass of the PBH:
$m_{\rm DM}\gtrsim 4.07\times 10^{-2}\left(\frac{M_{\rm in}}{1{\rm g}}\right)^{1/2}~{\rm GeV}.$ (32)
Considering the memory burden effect, we can utilize the final time for the semiclassical era, $t_{\rm q}$, as an effective lifetime of PBHs for particle production. In this case, we set $\tau=\tau^{0}(1-q^{3})$ instead of $\tau=\tau^{0}$ from Eq. (10). The WDM constraint with the memory burden effect is
$m_{\rm DM}\gtrsim 4.07\times 10^{-2}(1-q^{3})^{1/2}\left(\frac{M_{\rm in}}{1{\rm g}}\right)^{1/2}~{\rm GeV}.$ (33)
Thus, the allowed minimum DM mass is reduced by the memory burden effect.
Figure 1 shows the nature of Eq.(33). The left panel shows the lower limit of the WDM mass against the initial mass of the PBH. The left (right) gray region is excluded from the CMB (BBN) constraint. The blue solid line indicates the prediction without the memory burden effect, whereas the dotted lines show predictions with the memory burden effect at q=0.0050.9q=0.005-0.9. The right panel shows $q$ dependence on the lower limit of $m_{\chi}$. As expected, a large memory burden effect of the PBH reduces the WDM constraint.